Chris is buying supplies for a school fundraiser and has $56 to spend. He buys popcorn for $3 per bag and cotton candy for $7 per bag. He needs at least 7 bags of popcorn
Graph th boundary lines of the linear inequality

Answer by ikleyn(52781)

.
There are 2 (two) substantial inequalities in this problem.


Let X be the number of popcorn bags and Y be the number of cotton candy bags.


Then the total spending is  3X + 7Y dollars,  and it should be less or equal to 56 dollars.


So, your first inequality is


    3X + 7Y <= 56   dollars    (1)


The second inequality is


    X >= 7.                    (2)


The plot is below.



    


    Plot  3X + 7Y = 56 (red sloped line)  and  X = 7 (green vertical line)



The solution set to this inequality system is the TRIANGLE in QI below the red line and to the right of green line.



    There are also two "non-negativity" inequality  X >= 0  and  Y >= 0 to mention them for completeness.

    They say that the solution set is located in QI and is a part of QI.